identify the focus and the directrix of each parabola -5y^2 - x =0
Answer by lwsshak3(11628)

identify the focus and the directrix of each parabola
-5y^2 - x =0
5y^2=-x
y^2=-x/5
This is a parabola that opens leftwards
Its standard form of equation: ((y-k)^2=4p(x-h), (h,k)=(x,y) coordinates of vertex
For given parabola:
vertex:(0,0)
axis of symmetry: (x-axis) or y=0
4p=1/5
p=1/20
focus:(1/20,0) (p units from vertex on axis of symmetry)
Directrix: x=-1/20 (p units from vertex on axis of symmetry)